Solve or "x" Square root (3x+6) = 4
Answer by piglet162431(90)

First you need to inverse the square root,
Square 3x+6 ^2=3x+6. Now you've "canceled" the square.
But you have to do the same to 4
4^2= 16
Now. 3x+6 = 16
= 3x+6-6 = 16-6
3x = 10
3x/3 = 10/3
x = 10/3